A skier, starting from rest, skis down a slope from a height of 300 m to the horizontal ground level.

a) Show that the speed with which the skier reaches the ground level does

not depend on either the mass of the skier or the angle of incline of the slope. State any assumptions made.

b) Calculate the speed with which the skier reaches the ground level. 


1
Expert's answer
2021-11-14T17:11:42-0500

a)

by energy conservation law:

"mgh=\frac{mv^2}{2}"

the speed with which the skier reaches the ground level:

"v=\sqrt{2gh}"

where h is height


b)

"v=\sqrt{2\cdot9.8\cdot 300}=76.7" m/s